# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ VII-B Field Experiments and Pilot Projects
The Virginia Connected Corridors (VCC) project in the U.S. is led by the Virginia Department of Transportation and the Virginia Tech Transportation Institute [314]. It serves as a testbed for Connected and Autonomous Vehicle (CAV) applications. Since its launch in 2016, a fleet of 50 highly instrumented light vehicles and 60 Roadside Units (RSUs) have been deployed within the Northern Virginia Test Bed for application testing. Each vehicle is equipped to communicate via DSRC with the surrounding connected infrastructure. When the vehicles are outside the range of a DSRC-equipped RSU, they can use cellular technology for communication. A centralized cloud system is employed to manage the message traffic between connected vehicles. The vehicles can connect to the cloud through either DSRC or cellular networks. It has been noted that cellular communications have a longer latency, ranging from 1.5 to 3.5 seconds, compared to DSRC and are therefore used for applications where timely message receipt is not critical. The aim of this project is to provide an open development platform, allowing third-party developers to test and deploy applications that enhance roadway safety and mobility.
The iMOVE Cooperative Perception Project, led by the Australian Centre for Field Robotics at the University of Sydney in collaboration with Cohda Wireless, has demonstrated advancements in pedestrian detection around corners through the use of CPMs and Dedicated Short-Range Communication (DSRC) technology [315]. Notably, this project illustrates how CAVs can autonomously and safely interact with pedestrians, both walking and running, relying exclusively on CP information provided by an intelligent roadside unit (RSU) [315]. This project represents one of the pioneering demonstrations of urban vehicle automation that relies solely on CP data for safe and efficient operation in complex environments.
@@ VIII-A2 Real-world Communication Constraints
A promising approach is the integration of cross-layer models to assess how communication constraints affect CP algorithms [317, 318]. For example, combining network-layer routing with physical-layer channel modeling enables CP systems to anticipate the effects of channel fading or congestion on data transmission, facilitating adaptive adjustments in CP processes. Additionally, predictive perception models can enhance CP resilience by using historical data and probabilistic modeling to forecast surrounding vehicle behavior, maintaining situational awareness even during data loss. This reduces dependency on continuous communication, ensuring reliable CP performance in real-world conditions.
@@ VIII-B4 Collaborative AI Integration for Enhanced CP and V2X Communication
In the context of CP, each CAV must intelligently determine the information it requires from other connected vehicles and decide what information it should transmit to meet requests from other vehicles. This decision-making process is highly contingent on the dynamic traffic conditions. Therefore, each CAV’s AI system must efficiently manage CP’s necessary information, adapting to the current and near-future traffic environment. However, to optimize information flow management, each CAV must collectively consider the available communication resources at their disposal.
@@ VIII-B5 Migration of Advanced Perception Framework from Single Vehicle to Cooperative Vehicles
As the development of transformer model [326] in both computer vision and large language model areas, several advanced technologies have been proposed for single-vehicle perception and autonomous driving, e.g. end-to-end perception [327, 45], 3D occupancy grid prediction [328, 329], referring expression-based perception [330] with multi-modal large language model [331], self-supervised world-models for autonomous driving to boost the downstream tasks by using the features extracted from understanding and predicting/reconstruction of the environment [332, 333], and end-to-end autonomous driving [334, 335], etc. All of the aforementioned technologies have great potential to be extended into cooperative vehicle settings as well, with the expectation of enhancing the capability of autonomous driving. The most recent examples are [336, 337] and [338, 339, 340], in which the end-to-end autonomous driving framework and 3D occupancy grid prediction function are integrated with V2X cooperation, respectively, to achieve significant performance improvement on planning and intermediate perception outputs.

### v5

@@ VII-B Field Experiments and Pilot Projects
The Virginia Connected Corridors (VCC) project in the U.S. is led by the Virginia Department of Transportation and the Virginia Tech Transportation Institute [215]. It serves as a testbed for CAV applications. Since its launch in 2016, a fleet of 50 highly instrumented light vehicles and 60 Roadside Units (RSUs) have been deployed within the Northern Virginia Test Bed for application testing. Each vehicle is equipped to communicate via DSRC with the surrounding connected infrastructure. When the vehicles are outside the range of a DSRC-equipped RSU, they can use cellular technology for communication. A centralized cloud system is employed to manage the message traffic between connected vehicles. The vehicles can connect to the cloud through either DSRC or cellular networks. It has been noted that cellular communications have a longer latency, ranging from 1.5 to 3.5 seconds, compared to DSRC and are therefore used for applications where timely message receipt is not critical. The aim of this project is to provide an open development platform, allowing third-party developers to test and deploy applications that enhance roadway safety and mobility.
The iMOVE Cooperative Perception Project, led by the Australian Centre for Field Robotics at the University of Sydney in collaboration with Cohda Wireless, has demonstrated advancements in pedestrian detection around corners through the use of CPMs and Dedicated Short-Range Communication (DSRC) technology [216]. Notably, this project illustrates how CAVs can autonomously and safely interact with pedestrians, both walking and running, relying exclusively on CP information provided by an intelligent RSU [216]. This project represents one of the pioneering demonstrations of urban vehicle automation that relies solely on CP data for safe and efficient operation in complex environments.
@@ VIII-A2 Real-world Communication Constraints
A promising approach is the integration of cross-layer models to assess how communication constraints affect CP algorithms [218]. For example, combining network-layer routing with physical-layer channel modeling enables CP systems to anticipate the effects of channel fading or congestion on data transmission, facilitating adaptive adjustments in CP processes. Additionally, predictive perception models can enhance CP resilience by using historical data and probabilistic modeling to forecast surrounding vehicle behavior, maintaining situational awareness even during data loss. This reduces dependency on continuous communication, ensuring reliable CP performance in real-world conditions.
@@ VIII-B4 Collaborative AI Integration for Enhanced CP and V2X Communication
In the context of CP, each CAV must intelligently determine the information it requires from other connected vehicles and decide what information it should transmit to meet requests from other vehicles. This decision-making process is highly contingent on the dynamic traffic conditions. Therefore, each CAV’s AI system must efficiently manage CP’s necessary information, adapting to the current and near-future traffic environment. However, to optimize information flow management, each CAV must collectively consider the available communication resources at its disposal.
@@ VIII-B5 Migration of Advanced Perception Framework from Single Vehicle to Cooperative Vehicles
As the development of transformer models in both computer vision and large language model areas, several advanced technologies have been proposed for single-vehicle perception and autonomous driving, e.g., end-to-end perception [226], 3D occupancy grid prediction [227], referring expression-based perception [228] with multi-modal large language model [229], self-supervised world-models for autonomous driving to boost the downstream tasks by using the features extracted from understanding and predicting/reconstruction of the environment, and end-to-end driving. All of the aforementioned technologies have great potential to be extended into cooperative vehicle settings as well, with the expectation of enhancing the capability of autonomous driving. The most recent examples are [230, 231, 232, 233], in which the end-to-end autonomous driving framework and 3D occupancy grid prediction function are integrated with V2X cooperation, respectively, to achieve significant performance improvement on planning and intermediate perception outputs.